An armature has a resistance of 0.2 Ω at 150°C and the armature Cu loss is to be limited to 600 watts with a temperature rise of 55°C. If α0 for Cu is 0.0043/°C, what is the maximum current that can be passed through the armature?

1
Expert's answer
2021-12-07T20:14:40-0500

"R_T=R_0(1+\alpha_0\Delta\>T)"


"P_{max}=I^2_{max}\>R_T"


"T_0=150\u00b0c"

"R_0=0.2\Omega"

"\alpha_0=0.0043/°c"

"P_{max}=600w"

"\Delta\>T=55°c"

"R_T=0.2(1+0.0043\u00d755)"

"=0.2473"


"600=I^2_{max}\u00d70.2473"


"I^2_{max}=2426.2"

"I_{max}=49.26A"